OpenStudy (anonymous):

MOTHER CHECKING SONS HW HE PUT 8 AND 2 IS THIS RIGHT EXPLAIN AND ANWER PLS What is the solution to the system of equations? x=4y+6 x-3y=4

OpenStudy (anonymous):

u have x from the above equation,put it into the second one: \( (4y+6) - 3y = 4 \) so \(y+6 = 4\) so \( y=-2\) and u can put y in the above equation

OpenStudy (anonymous):

\(y=-2\) so \( x = 4(-2) + 6=-2\)